Most Jet fuel has an average molecular formula of C13H28 and a density of 804 g/L. Calculate the moles of carbon that are released from the fuel when a jet travels from Charlotte, NC to Los Angeles, CA, a distance of 3900 km. Assume the jet consumes 13L of fuel per Km


1
Expert's answer
2021-03-16T08:26:12-0400

Jet consumption of fuel = 13 L/km , Total distance travel= 3900 km

Total fuel consumption= "13\times 3900=50700" L,

Density of fuel = 804 g/L

Mass of fuel = "Volume\times density= 50700\times 804=40762800 g"

Molecular mass of fuel= "C_{13}H_{28}= 12\times 13 + 28\times 1=184 g/mole"

Total mole of "C_{13}H_{28}=\frac{40762800}{184}=221536.956" mole

Mole of C="13\times 221536.95=2879980.434" mole of C